OpenStudy (anonymous):

here is more i need help with if you divide and simplify. 4-2/4 4 the answers i have are a.4096 b.1 c.1/4096 d.16

OpenStudy (phi):

I think that is \( 4^4 \) in the denominator?

OpenStudy (phi):

people use the uparrow to show this 4^4 (don't type 4 4 because that is confusing)

OpenStudy (phi):

so is the problem \[ \frac{4^{-2}}{4^4} \] often written as 4^-2 / 4^4

OpenStudy (anonymous):

i sorry i don't know how to put it but down but the way phi put down was want i saying but i still need help with it

OpenStudy (radar):

\[4^{-2}\over4^{4}\]\[4^{(-2-4)}=4^{-6}=1/4^{6}\]

OpenStudy (radar):

\[4^{6}=4096\]

OpenStudy (radar):

The equation button beneath the left side of the data entry box can be helpful.

OpenStudy (anonymous):

now i got it i hope

OpenStudy (radar):

I hope so too, good luck.

OpenStudy (anonymous):

i still need help with this one if you multiply and simplify.3-2.3 there a small 6 next to 3

OpenStudy (phi):

is it \[ 3^{-2}\cdot 3^6 \]? type it as 3^-2 * 3^6 notice that both numbers have a 3 as the "base" the big number. which you multiply these numbers, keep the same base, but add the exponents (the small numbers to the upper right) so write down 3 and then put in the new exponent which is (can you tell me?)

OpenStudy (anonymous):

i don't know you are telling me to do this this new to every time i do this don't get the right answer i don't get the answer i have in my book i have a.1296 b.81 c.6561 d.1/81

OpenStudy (phi):

this is the problem 3^-2 * 3^6 can you pick out the exponents?

OpenStudy (phi):

the exponents are the little numbers in the upper right

OpenStudy (anonymous):

i still don't get it